# Hierarchical Multi-agent Reinforcement Learning for Cyber Network Defense

### v2

@@ 4.1. Hierarchical MARL Design
In the previous section, we primarily talked about our hierarchical approach at a high level from a single-agent perspective. In this subsection, we propose our method H-MARL Meta which adapts our hierarchical design for multi-agent tasks. H-MARL Meta is designed to overcome a multitude of difficulties induced by multi-agent training, such as environment and training instability caused by policy interdependencies. In particular, there are three key interdependencies that make learning in this setting difficult: (i) interdependence between each agent under a shared reward signal, (ii) between master and sub-policy performance, and (iii) between sub-policies under shared episodic returns.
@@
This first interdependence results in agents receiving rewards that are unrelated to actions they have taken. This is particularly challenging in our cyber defense setting as each agent interacts with disjoint sub-networks, but receives a shared reward considering the state of the whole network. This leads us to use IPPO De Witt et al. (2020) as the foundation of our approach, where each agent has a separate critic which only receives observations corresponding to their respective sub-network. Each agent is then trained in parallel along with its respective critic. This setup prevents the critics from being biased by occurrences outside their respective agent’s sub-network, resulting in greater stability and less bias in each agent.
@@
Secondly, the performance of the master policy depends on the performance of each sub-policy – poorly trained sub-policies can make otherwise optimal meta-actions sub-optimal – resulting in a biased master policy. To overcome this, we utilize a two-phase training approach seen in Algorithm 1 and Algorithm 2. Algorithm 1 describes how we train the sub-policies for each agent $i$ in parallel. We first define an expert policy $\pi_{E}$, which uses deterministic rules generated from domain knowledge to select meta-actions, along with a set of observation transformation functions $\{f_{c}\}_{c=1}^{k}$. At the start of each episode, we receive an initial observation from $\mathcal{M}$ and use it to initialize our history $h_{t}$. At each time step $\pi_{E}$ then uses $h_{t}$ to choose the best meta-action $A_{c}$. The sub-policy $\psi_{c}$ subsequently chooses a primitive action $a_{t}$ given its transformed history $f_{c}(h_{t})$. This action is used to sample the next observation and reward from $\mathcal{M}$ which is stored in the replay memory $\mathcal{D}_{c}$ of $\psi_{c}$. Each policy is then updated with PPO Schulman et al. (2017) on its respective replay memory $\mathcal{D}_{c}$. This design allows the sub-policies to address only their related tasks and train to near-optimal while avoiding instability caused by a trained master policy. Additionally, this allows us to solve our third form of interdependence, as the deterministic, static expert policy $\pi_{E}$ allows for stability in the training of each sub-policy.
Algorithm 1 Sub-policy training for agent $i$
@@
2:Sub-Policies $\{\psi_{c}\}_{c=1}^{k}$, Replay Memories $\{\mathcal{D}_{c}\}_{c=1}^{k}$, Transformations $\{f_{c}\}_{c=1}^{k}$, Episode Length $T$, Iterations $N$
@@
Algorithm 2 Master policy training for agent i
1:Dec-POMDP $\mathcal{M}$, Transformations $\{f_{c}\}_{c=1}^{k}$
2:Master Policy $\pi_{m}$, Sub-Policies $\{\psi_{c}\}_{c=1}^{sp}$, Replay Memory $\mathcal{D}_{m}$, Episode Length $T$, Iterations $N$
@@
Algorithm 2 provides the second phase of training for the master policy. Here sub-policies $\{\psi\}_{1}^{k}$ trained with Algorithm 1 are kept frozen and just used to generate primitive actions. We define a master policy $\pi_{m}$ whose action space is the number of sub-policies $k$. Similar to before, at each time step we sample an action $c_{t}\sim\pi_{m}(h_{t})$ given updated history $h_{t}$. The sub-policy $\psi_{c_{t}}$ is invoked to sample action $a_{t}$ to take a step in $\mathcal{M}$ given transformed history $f_{c}(h_{t})$. The current history $h_{t}$, action $c_{t}$ and reward from $\mathcal{M}$ given action $a_{t}$ are then stored in the replay memory $\mathcal{D}_{m}$. Finally, the master policy $\pi_{m}$ is updated with PPO given trajectories stored in $\mathcal{D}_{m}$.
Instantiating H-MARL in cybersecurity. Figure 2 illustrates the sub-task partition in a cyber environment such as CybORG CAGE 4. We identify three types of sub-tasks: investigate host, recover host, and control traffic between zones. The state-level abstraction used to partition the tasks refers to the presence of indicators of compromise within an agent’s security zone(s). Thus, if IOCs such as malicious files are identified on a host, the agent will choose to recover, either by removing the malware or by restoring the machine to a clean state. Furthermore, if network IOCs are detected, such as incoming scanning activity captured by decoy services, then the sub-policy handling traffic control can choose to block traffic originating from the compromised subnet. Each of these sub-policies will receive the relevant subspace of the observation space and the subset of relevant actions.
Figure 2. Instantiating H-MARL in cybersecurity. The master policy uses state abstractions related to the presence of IOCs. Each sub-policy learns a specific behavior based on its own subset of actions and observation space.
In the CAGE 4 setup, blocking the traffic between neighboring subnets is not particularly useful for two reasons: (i) it prevents green agents from accessing remote services, which penalizes the blue team, and (ii) red agents are active mostly within their own assigned subnet, and rarely reach remotely into other subnets, a setup that ensures scaling with large networks. In the rest of the paper, unless otherwise specified, we discuss defenses that do not use block or allow traffic actions, focusing on the Investigate and Recover sub-policies. We explore traffic control guided by communication among blue agents in supplemental material Section A.4.
4.2. Observation Space Design
@@
Figure 3. Observation space components. The basic CybORG observation is enhanced with indicators of compromise.
@@
Enhancing observation with IOCs. Indicators of compromise (IOCs) Asiri et al. (2023) are signs or evidence of a cyber threat being present in the network. IOCs are generally classified in three categories: atomic (IP addresses, malware names, registry keys, process names, URLs, etc.), computed (hash of a malicious file), and behavioral Hutchins et al. (2011). MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) Strom et al. (2018) is a comprehensive database of adversarial behaviors observed in real-world attacks.
@@
Observation space evaluation. Figure 4 shows the contribution of each of our enhancements to the observation space design for a blue agent trained with a decentralized actor-critic PPO architecture, using the same hyperparameters from Section 5. Compared to the basic CybORG observation, keeping track of history on suspicious events offers a small performance boost (12% increase in reward). The biggest gain, however, comes from incorporating indicators of compromise related to malicious files (an additional increase in reward of 42%). These files are detected when blue agents perform Analyse actions on hosts in their assigned subnet. Access to decoys, another clear indicator of adversarial behavior, further improves the defense strategy of the blue team by 11%.
Figure 4. Blue team reward for different observation space designs. We incrementally add each of our enhancements to the observation space, to show their individual contribution. Incorporating history and IOCs provides a high performance boost. (MARL Decentralized training)
5. Experimental Evaluation
@@
Training configuration. Our experiments use the state-of-the-art actor-critic PPO algorithm Schulman et al. (2017). The actor and critic are represented by two feedforward neural networks with two hidden layers and 256 neurons per layer. The training hyperparameters have been tuned to the following values: a learning rate of 0.00005, a 0.99 discount factor, and a train buffer size of 1 million samples. The SGD algorithm uses the Generalized Advantage Estimation (GAE) function, a minibatch size of 32,768 within each epoch, with 30 SGD iterations in each outer loop. Evaluation results are averaged across 100 randomised episodes, where each episode is 500 time steps long, and are accompanied by standard deviation information.
@@
Duration of actions. For increased realism actions in the environment take a specified number of steps to execute. Some of the longer actions are Exploit Remote Services (4 steps) and Restore host (5 steps). In our implementation, while an action is in progress in the environment, we associate the “in-progress” state with a special observation and mask out other actions except sleep to guide PPO training.
5.1. H-MARL Performance
@@
MARL Baselines. We compare the hierarchical architecture with two single-policy MARL paradigms Lyu et al. (2021): Decentralized Training Decentralized Execution (DTDE), and Centralized Training Decentralized Execution (CTDE). The DTDE baseline uses an actor-critic architecture that learns a decentralized policy and critic for each of the agents locally (MARL Decentralized). The CTDE baseline uses the centralized critic approach presented in MADDPG Lowe et al. (2017). Since each learning blue agent is required to guide its strategy based on the joint team reward, we augment the critic with state and actions of all blue agents on the team, while the actor only has access to local observations. Note that our MARL Centralized Critic baseline uses the global state instead of incomplete agent observations to compute the joint value function. This is a reasonable assumption during training, in an effort to provide an unbiased and up-to-date critic for a strong baseline Lyu et al. (2021).
@@
H-MARL Methods. We evaluate two hierarchical methods, H-MARL Expert and H-MARL Meta. H-MARL Expert implements Algorithm 1, where the master policy is replaced with a rule informed by domain knowledge: If IOCs are detected, Recover; otherwise, Investigate. The goal is to recover right away to minimize the attacker’s damage. H-MARL Meta implements Algorithms 1 and 2, following a curriculum style approach Bengio et al. (2009): the Recover and Investigate sub-policies pre-trained until convergence, are maintained fixed while training the master policy. H-MARL Collective is an additional baseline that attempts to learn both the master and sub-policies from scratch, simultaneously. All the hierarchical variants use IPPO, in the decentralized actor-critic framework.
Figure 5. Average training return for all algorithms. H-MARL Expert is guided by a rule-based master policy and performs best. H-MARL Meta converges 35×3-5$\times$3 - 5 × faster than MARL Decentralized.
@@
The training results for the blue agents protecting against the default red agent in CAGE 4 are presented in Figure 5. The H-MARL Collective method performs the worst, in line with previous work Frans et al. (2017) that also observed the sub-optimal performance of updating both sub-policies and the master policy at the same time. Both single-policy methods MARL Decentralized and MARL Centralized Critic converge to a high reward, with a clear advantage for the shared critic method that estimates the return based on joint information.
@@
As expected, the H-MARL Expert performs best ($-129.53$ reward), given that recovery actions are carried out promptly, before the attack amplifies. H-MARL Meta reaches a similar reward to MARL Decentralized ($-181.62$). However, training a master policy is significantly faster than training a single policy from scratch (about 3-5 times faster), as we are only tasked with choosing the correct sub-policy, rather than also learning the primitive sub-tasks, i.e., how to recover or investigate hosts. H-MARL Meta invokes the fixed sub-policies that have been pre-trained with H-MARL Expert to take a step though the environment. The ability to learn how to combine sub-tasks to solve higher-level tasks effectively with H-MARL Meta is particularly important in situations where defining an expert rule is difficult. We discuss such a scenario in Section 5.5.
5.2. Evaluation against different adversaries
@@
Table 2. Mean reward against different adversaries, under different learning frameworks. Evaluation is carried out after training has converged. Results are averaged over 100 evaluation episodes, with each episode consisting of 500 time steps.
Opponent MARL MARL H-MARL H-MARL Expert H-MARL Meta H-MARL Meta
Decentralized Centralized Critic Collective with fine-tuning
Default Red $-179.8\pm 92.98$ $-245.66\pm 132.98$ $-237.61\pm 102.6$ $\mathbf{-129.53\pm 44.60}$ $-181.62\pm 65.85$ $-186.24\pm 81.28$
Stealthy Red $-165.8\pm 53.45$ $-217.52\pm 115.06$ $-204.18\pm 83.24$ $\mathbf{-99.54\pm 38.28}$ $-184.76\pm 91.40$ $-162.51\pm 57.97$
Aggressive Red $-227.93\pm 87.98$ $-255.98\pm 138.70$ $-282.5\pm 120.50$ $\mathbf{-118.16\pm 37.74}$ $-207.66\pm 86.48$ n/a
Impact Red $-247.15\pm 71.53$ $-332.96\pm 108.13$ $-350.21\pm 115.25$ $\mathbf{-173.17\pm 64.04}$ $-278.78\pm 97.76$ $-264.96\pm 81.53$
@@
We evaluated so far the performance of blue agents against the default red agent in CAGE 4, but we are interested in how the policy generalizes against other red attacks. We simulate red agent attackers that vary in their ability to circumvent the defense by employing more aggressive scanning, stealthier behavior, or a stronger focus on impacting critical services.
@@
In the CybORG environment, red agents use finite state machine transitions to determine what actions to take for each known host. For a detailed description of the default finite state red agent and the transition matrix please reference CybORG ([n.d.]). We consider four different finite state adversaries: (1) Default Red – equal choice between the two available service discovery actions (stealthy and aggressive), and equal split between the two attack objectives (Impact and Degrade Service); (2) Aggressive Red – always performs aggressive service discovery, a short duration action (1 time step) that has a high chance of being detected by blue agents (0.75); (3) Stealthy Red – stealthy service discovery, characterized by low detection rate (0.25) and long duration (3 time steps); (4) Impact Red – fully committed to impacting the critical OT service, without attempting to degrade other services.
@@
The evaluation results against these four different adversaries are presented in Table 2. Impact Red is the strongest attacker collecting the highest reward against the blue team, in this zero-sum cyber game. Successful impact actions receive the highest reward of -10 during the second and third mission phases, more than any other network compromise. For comparison, users’ failed access to degraded services only costs the blue team between -1 and -3. Aggressive Red is the second strongest opponent: the higher chance of being detected by blue agents is offset by the short duration of service discovery actions (1 time step), which enable this attacker to carry out more frequent attempts. Conversely, and somewhat non-intuitively, Stealthy Red poses less of a threat to blue agents due to the longer duration spent in covert scanning, before attempting to exploit vulnerabilities.
@@
H-MARL Expert is the best strategy against all four red agents and H-MARL Meta attains similar reward to the single policy baseline method (MARL Decentralized), but training converges significantly faster, as shown in the previous section. We confirm that H-MARL Meta performs better than H-MARL Collective across all adversaries. Interestingly, it is also more effective than the MARL Centralized Critic method. Even though the centralized critic was trained jointly, using the global state, it is unable to utilize the shared information effectively on the individual actors during evaluation. This limitation of centralized critic policies has been previously investigated in the literature Lyu et al. (2021).
5.3. H-MARL Transferability

### v3

@@ 4.1 Hierarchical MARL Design
We now propose our methods H-MARL Expert and H-MARL Meta, which are designed to overcome a multitude of challenges induced by multi-agent training, such as environment and training instability. In particular, there are three key inter-dependencies that make learning in this setting difficult: (i) interdependence between each agent under a shared reward signal, (ii) between master and sub-policy performance, and (iii) between sub-policies under shared episodic returns.
@@
The first interdependence results in agents receiving rewards that are not related to actions they have taken. This is particularly challenging in our cyber defense setting as each agent interacts with disjoint sub-networks, but receives a shared reward considering the state of the whole network. This leads us to use IPPO (de2020independent, ) as the foundation of our approach, where each agent has a separate critic which only receives observations corresponding to their respective sub-network. Each agent is then trained in parallel along with its respective critic. This setup prevents the critics from being biased by occurrences outside their respective agent’s sub-network, resulting in greater stability and less bias in each agent. Secondly, the performance of the master policy depends on the performance of each sub-policy – poorly trained sub-policies can make otherwise optimal meta-actions sub-optimal – resulting in a biased master policy. To overcome this, we utilize a two-phase training approach seen in Algorithm 1 and Algorithm 2. For both algorithms, the training is guided by the reward signal received from the environment (see Section 3).
Algorithm 1: Sub-policy training (H-MARL Expert). Our first method, H-MARL Expert, uses an expert master policy $\pi_{E}$ defined by domain expertise and only trains sub-policies for each agent, using Algorithm 1. The H-MARL Expert Pipeline is presented in Figure S-2 from the Supplemental Materials.
@@
At the start of each episode, we receive an initial observation from $\mathcal{M}$ and use it to initialize our history $h_{t}$. At each time step, $\pi_{E}$ then uses $h_{t}$ to choose the best meta-action $A_{c}$. Next, the sub-policy $\psi_{c}$ (corresponding to meta-action $A_{c}$) chooses a primitive action $a_{t}$ given its transformed history $f_{c}(h_{t})$. This action is used to sample the next observation and reward from $\mathcal{M}$ which is stored in the replay memory $\mathcal{D}_{c}$ of $\psi_{c}$. Each policy is then updated with PPO on its respective replay memory $\mathcal{D}_{c}$. This design allows the sub-policies to address only their related tasks and train to near-optimal while avoiding instability caused by a trained master policy. Additionally, this allows us to solve our third form of interdependence, as the deterministic, static expert policy $\pi_{E}$ allows for stability in the training of each sub-policy.
Algorithm 1 Sub-policy training (H-MARL Expert)
@@
2:Sub-Policies $\{\psi_{c}\}_{c=1}^{k}$, Replay Memories $\{\mathcal{D}_{c}\}_{c=1}^{k}$, Transformations $\{f_{c}\}_{c=1}^{k}$, Episode Length $T$, Iterations NN
@@
Algorithm 2 Master policy training (H-MARL Meta)
1:Dec-POMDP $\mathcal{M}$, Sub-Policies $\{\psi_{c}\}_{c=1}^{sp}$, Transformations $\{f_{c}\}_{c=1}^{k}$
2:Master Policy $\pi_{m}$, Replay Memory $\mathcal{D}_{m}$, Episode Length $T$, Iterations NN
@@
H-MARL Expert in cybersecurity. Figure 2 illustrates the Expert master policy $\pi_{E}$ for partitioning the sub-tasks in CybORG CAGE-4. We identify three types of sub-tasks: investigate host, recover host, and control traffic between zones. The state-level abstraction used to partition the tasks refers to the presence of indicators of compromise (IOCs) within an agent’s security zone(s). This partitioning is defined via Expert Rules, including: (1) If IOCs (malicious files) are detected on a host, the agent will choose the Recover subpolicy, which selects either to remove the malware or to restore the machine to a clean state; (2) If network IOCs are detected, then the Control Traffic subpolicy is chosen; (3) Otherwise, the agent will Investigate.
Figure 2: H-MARL in cybersecurity. The Expert Master Policy knows this IOC-based partitioning, while the Meta Master Policy is learning it using frozen subpolicies.
Algorithm 2: Master policy training (H-MARL Meta). Algorithm 2 provides the second phase of training for the master policy. Here sub-policies $\{\psi\}_{1}^{k}$ trained with Algorithm 1 are kept frozen and just used to generate primitive actions. The H-MARL Meta Pipeline is presented in Figure S-3 from the Supplemental Materials.
We define a master policy $\pi_{m}$ called H-MARL Meta whose action space size is the number of sub-policies $k$. Similar to before, at each time step we sample an action $c_{t}\sim\pi_{m}(h_{t})$ given updated history $h_{t}$. The sub-policy $\psi_{c_{t}}$ is invoked to sample action $a_{t}$ to take a step in $\mathcal{M}$ given transformed history $f_{c}(h_{t})$. The current history $h_{t}$, action $c_{t}$ and reward from $\mathcal{M}$ given action $a_{t}$ are then stored in the replay memory $\mathcal{D}_{m}$. Finally, the master policy $\pi_{m}$ is updated with PPO given trajectories stored in $\mathcal{D}_{m}$.
4.2 Observation Space Design
@@
Figure 3: Observation space components. The basic CybORG observation is enhanced with IOCs.
@@
Enhancing observation with IOCs. Indicators of compromise (IOCs) (asiri2023, ) are signs or evidence of a cyber threat being present in the network. IOCs are generally classified in three categories: atomic (IP addresses, malware names, registry keys, process names, URLs, etc.), computed (hash of a malicious file), and behavioral (hutchins2011intelligence, ). MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) (strom2018mitre, ) is a comprehensive database of adversarial behaviors observed in real-world attacks.
@@
Observation space evaluation. Figure 4 shows the contribution of each of our enhancements to the observation space design for a blue agent trained with a decentralized actor-critic PPO architecture, using the same hyper-parameters from Section 5. Compared to the basic CybORG observation, keeping track of history on suspicious events offers a small performance boost (12% increase in reward). The biggest gain, however, comes from incorporating indicators of compromise related to malicious files (an additional increase in reward of 42%). These files are detected when blue agents perform Analyse actions on hosts in their assigned subnet. Access to decoys, another clear indicator of adversarial behavior, further improves the defense strategy of the blue team by 11%. Note that the H-MARL architecture can not be applied on the basic CybORG observation space. This is because task partitioning in H-MARL is based on indicators of compromise, which are not tracked in the original CybORG. In the rest of the paper, we consider the enhanced observation space with history, IOCs, and decoys for training all agents.
Figure 4: Blue team reward for different observation space designs. We incrementally add each of our enhancements to the observation space, to show their individual contribution. Incorporating history and IOCs provides a high performance boost. (MARL Decentralized training)
5 Experimental Evaluation
@@
Training configuration. Our experiments use the state-of-the-art actor-critic PPO algorithm (schulman2017proximal, ). The actor and critic are represented by two feedforward neural networks with two hidden layers and 256 neurons per layer. The training hyper-parameters have been tuned to the following values: a learning rate of $5\times 10^{-5}$, a discount factor of 0.99, and a train buffer size of 1 million samples. The SGD algorithm uses the Generalized Advantage Estimation (GAE) function, a mini-batch size of 32,768 within each epoch, with 30 SGD iterations in each outer loop. Evaluation results are averaged across 100 randomized episodes, where each episode is 500 time steps long, and are accompanied by standard deviation information. The network topology in the CybORG environment is randomized and the models are trained on topologies with varying configurations, which ensures that they generalize to different network environments. The size of the network at initialization is randomly chosen between 32 and 128 hosts across 8 subnets, and up to 10 services are selected randomly to be placed on each host. All our models have been trained on 30 versions of the network, with separate workers collecting the experiences in their own network.
@@
Duration of actions. For increased realism, actions take more than one step to execute. Some of the longer actions are Exploit Remote Services (4 steps) and Restore host (5 steps). In our implementation, while an action is in progress in the environment, we associate the “in-progress” state with a special observation and mask out other actions except sleep to guide PPO training.
5.1 H-MARL Performance
@@
MARL Baselines. We compare the hierarchical architecture with two single-policy MARL paradigms (lyu2021contrastingcentralizeddecentralizedcritics, ): Decentralized Training Decentralized Execution (DTDE), and Centralized Training Decentralized Execution (CTDE). The DTDE baseline uses an actor-critic architecture that learns a decentralized policy and critic for each of the agents locally (MARL Decentralized). The CTDE baseline uses the centralized critic approach presented in MAPPO (mappo, ). Since each learning blue agent is required to guide its strategy based on the joint team reward, we augment the critic with state and actions of all blue agents on the team, while the actor only has access to local observations. Note that our MARL Centralized Critic baseline uses the global state instead of incomplete agent observations to compute the joint value function. This is a reasonable assumption during training, in an effort to provide an unbiased and up-to-date critic for a strong baseline (lyu2021contrastingcentralizeddecentralizedcritics, ).
@@
H-MARL Methods. We evaluate two hierarchical methods, H-MARL Expert and H-MARL Meta. H-MARL Expert implements Algorithm 1, where the master policy is replaced with a rule informed by domain knowledge: If IOCs are detected, Recover; otherwise, Investigate. The goal is to recover right away to minimize the attacker’s damage. H-MARL Meta implements Algorithms 1 and 2, following a curriculum style approach (Bengio2009, ): the Recover and Investigate sub-policies pre-trained until convergence, are maintained fixed while training the master policy. H-MARL Collective is an additional baseline that attempts to learn both the master and sub-policies from scratch, simultaneously. All the hierarchical variants use IPPO, in the decentralized actor-critic framework.
@@
The training process for the blue agents is presented in Figure 5. The H-MARL Collective method performs the worst, in line with previous work (Frans2018, ) that also observed the sub-optimal performance of updating both sub-policies and the master policy at the same time. Both single-policy methods MARL Decentralized and MARL Centralized Critic converge to a high reward, with a clear advantage for the shared critic method that estimates the return based on joint information.
Figure 5: Average training return for all algorithms. H-MARL Expert is guided by a rule-based master policy and performs best. H-MARL Meta converges $3-5\times$ faster than MARL Decentralized.
@@
As expected, the H-MARL Expert performs best (129.53-129.53 reward), given that recovery actions are carried out promptly, before the attack amplifies. H-MARL Meta reaches a similar reward to MARL Decentralized (181.62-181.62). However, training a master policy is significantly faster than training a single policy from scratch (about 3-5 times faster), as we are only tasked with choosing the correct sub-policy, rather than also learning the primitive sub-tasks, i.e., how to recover or investigate hosts. H-MARL Meta invokes the fixed sub-policies that have been pre-trained with H-MARL Expert to take a step though the environment. The ability to learn how to combine sub-tasks to solve higher-level tasks effectively with H-MARL Meta is particularly important in situations where defining an expert policy is difficult. We discuss such a scenario in Section 5.4.
5.2 Evaluation against different adversaries
@@
Table 2: Mean evaluation reward against different adversaries, under different learning frameworks.
Opponent Default Red Stealthy Red Aggressive Red Impact Red
MARL Decentralized $179.8\pm 92.98$ $165.8\pm 53.45$ $227.93\pm 87.98$ $247.15\pm 71.53$
MARL Centralized Critic $245.66\pm 132.98$ $217.52\pm 115.06$ $255.98\pm 138.70$ $332.96\pm 108.13$
H-MARL Collective $237.61\pm 102.6$ $204.18\pm 83.24$ $282.5\pm 120.50$ $350.21\pm 115.25$
H-MARL Expert $\mathbf{-129.53\pm 44.60}$ $\mathbf{-99.54\pm 38.28}$ $\mathbf{-118.16\pm 37.74}$ $\mathbf{-173.17\pm 64.04}$
H-MARL Meta $181.62\pm 65.85$ $184.76\pm 91.40$ $207.66\pm 86.48 &$ $278.78\pm 97.76$
H-MARL Meta with fine-tuning $186.24\pm 81.28$ $162.51\pm 57.97$ n/a $264.96\pm 81.53$
@@
We evaluated so far the performance of blue agents against the default red agent in CAGE 4, but we are interested in how the policy generalizes against other red attacks. We simulate red agent attackers that vary in their ability to circumvent the defense by employing more aggressive scanning, stealthier behavior, or a stronger focus on impacting critical services. In the CybORG environment, red agents use finite state machine transitions to determine what actions to take for each known host. For a detailed description of the default finite state red agent and the transition matrix please see cyborg_finite_state_red .
@@
We consider four different finite state adversaries: (1) Default Red – equal choice between the two available service discovery actions (stealthy and aggressive), and equal split between the two attack objectives (Impact and Degrade Service); (2) Aggressive Red – always performs aggressive service discovery, a short duration action (1 time step) that has a high chance of being detected by blue agents (0.75); (3) Stealthy Red – stealthy service discovery, characterized by low detection rate (0.25) and long duration (3 time steps); (4) Impact Red – fully committed to impacting the critical OT service, without attempting to degrade other services.
@@
The evaluation results against these four different adversaries are presented in Table 2. Impact Red is the strongest attacker collecting the highest reward against the blue team, in this zero-sum cyber game. Successful impact actions receive the highest reward of -10 during the second and third mission phases, more than any other network compromise. For comparison, users’ failed access to degraded services only costs the blue team between -1 and -3. Aggressive Red is the second strongest opponent: the higher chance of being detected by blue agents is offset by the short duration of service discovery actions (1 time step), which enable this attacker to carry out more frequent attempts. Conversely, and somewhat non-intuitively, Stealthy Red poses less of a threat to blue agents due to the longer duration spent in covert scanning, before attempting to exploit.
@@
H-MARL Expert is the best strategy against all four red agents and H-MARL Meta attains similar reward to the single policy baseline method (MARL Decentralized), but training converges significantly faster, as shown in the previous section. We confirm that H-MARL Meta performs better than H-MARL Collective across all adversaries. Interestingly, it is also more effective than the MARL Centralized Critic method. Even though the centralized critic was trained jointly, using the global state, it is unable to utilize the shared information effectively on the individual actors during evaluation. This limitation of centralized critic policies has been previously investigated in the literature lyu2021contrastingcentralizeddecentralizedcritics .
5.3 H-MARL Transferability